OpenStudy (arbyj):

"A line passes through the points (-6, 3) and (5, -2)" how do i make this into an equation?

OpenStudy (anonymous):

do you know how to find the slope?

OpenStudy (arbyj):

Yes, but I'm not sure how to make into an equation

OpenStudy (anonymous):

think about point slope form once you get the slope. and if you think about it. point slope form and the slope formula are the same thing reconfigured.

OpenStudy (anonymous):

once you get slope. you just have to use 1 set of points to get the equation.

OpenStudy (kainui):

You know you're going to end up with y=mx+b right? So if you find the slope, then you can find b since you already have x,y, and m. Plug and chug!

OpenStudy (arbyj):

So I can use any set of the numbers listed to plug into y=mx+b?

OpenStudy (arbyj):

@lonnie455rich

OpenStudy (anonymous):

well. you can do that. but here is my thought process to add to his. given the information you provided \[m=\frac{ y-y }{ x-x }\] to get the slope \[y-3=m(x+6)\] then solve for y and you get \[y=mx+b\]